OpenStudy (anonymous):

y2+5y divided by y ______ _____ y^2+4y-5 y+1

OpenStudy (anonymous):

what is the question lol

OpenStudy (anonymous):

\[\frac{y^2+5y}{y^2+4y-5}\times \frac{y+1}{y}\] is a start. then factor and cancel what you can

OpenStudy (anonymous):

\[\frac{y(y+5)(y+1)}{(y+5)(y-1)y}\] \[\frac{y+1}{y-1}\]

OpenStudy (anonymous):

Multiply the numerator by the reciprocal of the denominator!